1.3 Convert to base 6: 3BA.25y, (do all of the arithmetic in decimal). 1.4 (a) Convert to hexadecimal: 1457.11 1o Round to two digits past the hexadecimal point. (b) Convert your answer to binary, and then to octal. (c) Devise a scheme for converting hexadecimal directly to base 4 and convert your answer to base 4. (d) Convert to decimal: DEC.Ar

help solving
Answer #1

3BA.25 base 14
= 3*14^2 + 11*14 + 12 + 2/14 + 5/(14^2) = 752.16836734 base 10

Decimal part

6|752
------
125 - 2
20 - 5
3 - 2
Go from bottom to up = 3252

Fraction part

.16836734 * 6 = 1.01020404
0.01020404 * 6 = 0.01020404
0.01020404 * 6 = 0.36734544
0.36734544 * 6 = 2.20407264
0.20407264 * 6 = 1.22443584
0.22443584 * 6 = 1.34661504
.34661504 * 6 = 2.07969024
.07969024 * 6 = 0.47814144
Come from top to bottom = 10021120

= 3252.10021120 base 6

Note: One question at a time please -- Policy of Chegg